Parth (parthkohli):

How can I perform operations on exponents and powers? See below to see the solution.

Parth (parthkohli):

Exponents in fractions: MEDALS WOULD BE APPRECIATED. I will explain the basic properties of exponents and powers by using variables a,b, m and n. Basic structure of exponents: In the expression \[2^{3}\], 2 is the base and 3 is the exponent.This means that 2 is multiplied by itself 3 times. So, \[2^{3} = 2*2*2 = 8\]. Multiplying with the same base: \[\LARGE a^{n} * a^{m} = a^{n+m}\] Multiplying with the same exponent but different base: \[\LARGE a^{n} * b{n} = {(a + b)^{n}} \] Dividing with the same base: \[\LARGE {{a^{m}} \over {a^{n}} = {a}^{m-n}}\] Multiplying an exponent by an exponent: \[\LARGE {(a^{m})^{n}} = {a ^ {mn}} \] Note that we cannot simply get the value of: \[ \LARGE {a^{m}} - {a^{n}} \] or \[ {a^{m}} + {a^{n}} \] Exponents in negative: \[\LARGE {a ^{-m}} = {1 \over a^{m}}\] Exponents in fractions: \[\LARGE {a ^ {1 \over m}} = \sqrt[m]{a}\]

OpenStudy (anonymous):

\[a^{n}*bn=(a+b)^{n}\] ? please explain

Parth (parthkohli):

I meant a^n * b^n

Parth (parthkohli):

Also, dividing by the same base has an error.

OpenStudy (anonymous):

wouldn't it be a^n+nab+b^n

OpenStudy (anonymous):

follow the laws of exponent

OpenStudy (lgbasallote):

i think i disagree that \(\large a^n \times b^n = (a+b)^n\) assuming n = 2... \((a+b)^2 = a^2 + 2ab + b^2 \ne (a^2 \times b^2)\)

OpenStudy (lgbasallote):

but anyways good job :)

Parth (parthkohli):

Yes, I forgot to put that one in. I also tried that last year.

OpenStudy (anonymous):

yes, PK is doing good stuff here

Parth (parthkohli):

\[\LARGE {a^{m} \over a^{n}} = {a^{m - n}} \]

OpenStudy (lgbasallote):

i also disagree that \(\Large (a+b)^n = a^n + nab + b^n\) lol...it only works in square.. for example \((a+b)^3 \ne a^3 + 3ab + b^3\) since \((a+b)^3 = (a^2 + 2ab + b^2)(a+b)\). just clearing the misconceptions :D so there is no rule for \(\large a^n \times b^n\)

Parth (parthkohli):

Another pattern that I noticed is that if a<m and a and m are more 2, \[\LARGE a^{m} > m^{a}\]

OpenStudy (anonymous):

good job

OpenStudy (lgbasallote):

well there is..it's simply \(a^n b^n\)

Parth (parthkohli):

I suck at tutorials lol

OpenStudy (lgbasallote):

nahh it's fine :DDD you'll get there hehe

OpenStudy (lgbasallote):

remember that these corrections are not critics..they are merely for references

Parth (parthkohli):

I could have easily written this on paper, but in latex there are lotta errors :P... Thanks lg. I hope you like my badly written yet fair tutorials.

OpenStudy (lgbasallote):

you might wanna add this: \(\LARGE a^{\frac{n}{m}} = \sqrt[m]{a^n}\)

Parth (parthkohli):

See the last one bro.. there it is.

OpenStudy (lgbasallote):

nahh..you only put 1/m :P hehe

Parth (parthkohli):

Oh yeah. :D

OpenStudy (lgbasallote):

and yes i do like it :) i couldnt have listed all of those down myself..i have poor memory ahah..this can help a lot of people so keep bumping it

Parth (parthkohli):

=)

OpenStudy (callisto):

\[a^n \times b^n \neq (a+b)^n\] But \[a^n \times b^n = (ab)^n\] I think only :)

Parth (parthkohli):

Haha the latex confused me so much that I got the math wrong. Thanks for correcting me :)

OpenStudy (callisto):

You're welcome